OpenStudy (anonymous):

How do you figure out if there is a triangle, or no triangle? I'm getting vex... For example: a=10, b=16.6, A=27 degree (produces two triangles). Example 2: a=18, b=25, A=70 degree (No Triangle). Please explain, why.

OpenStudy (anonymous):

I did notice that these two problems are quiet different.

OpenStudy (turingtest):

I'm thinking that one way to see it is with the law of sines...

OpenStudy (turingtest):

\[{a \over \sin A}={b \over \sin B}\to \sin B={b \over a} \sin A={25\over18}\sin(70)\approx1.3\]but trig functions can never be greater than one, so \[B=\sin^{-1}(1.3)\]is undefined, therefor no triangle.

OpenStudy (turingtest):

for the first triangle \[\sin B={16.6\over10}\sin(27)\approx0.75<1\]so angle B can be calculated, hence this obeys the law of sines and is a triangle.

OpenStudy (anonymous):

Thank you much.